Balcom Enterprises is planning to introduce a new product that will sell for $150 a unit. Manufacturing cost estimates for 25,500 units for the first year of production are:

  • Direct materials $1,530,000
  • Direct labor $570,000 (based on $19 per hour × 30,000 hours)

Although overhead has not been estimated for the new product, monthly data for Balcom's total production for the last two years has been analyzed using simple linear regression. The analysis results are as follows:

Dependent variable Factory overhead costs
Independent variable Direct labor hours
Intercept $ 131,000
Coefficient on independent variable $ 6.00
Coefficient of correlation 0.944
R2 0.836


Based on this information, what is the total overhead cost for an estimated activity level of 40,000 direct labor-hours?

$786,000

$141,000

$371,000

$240,000
Answer #1

Answer : $371,000

Cost Equation = Intercept +( Direct labor hours *Co-efficient independent variable )

Cost at 40,000 Direct labor hours

= 131,000+(6.00*40,000) = 131,000+240,000 = 371,000 (Answer)